LUCIFER Com2
UT 2008 December 09

Observers: N.Ageorges, W.Seifert, M.Juette, F. Eisenhauer. J. Hill (@LBT)
Telescope Operator: A. Ceranski
SW Support: T. Edgin on call

Summary:

It almost cleared near sunset, but we never got rid of the fog. Then another front with wind and snow moved through. The whole night has been lost to weather.

Details:

-- JohnHill - 09 Dec 2008
Topic revision: r2 - 09 Dec 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback